Find the Taylor series around a = 2 for h(x) = 72 ret

AND write the series into a summation notation formula.
Answer #1

A Taylor serise is given by In our case, f 60 = P(x) = * FK 962-90 R=0 R1 lan (x-gir R= 0 flo 6 = f (x = 3 - 2 xti Evaluate t- +32 ( +382 (0+ 12+2) (can be seen here Evaluate the Ish la given Point derivative at the (f/2) = 10 find the and delivative18 84 +4x3 + 6x² + 4x+1 given Point (f. (2)) = 2 find the 9th deh vative: gi (4) 66 7 = ((37 (60) - Go4s +bar2 14.673) 72 XS=__ 366 x2 + 6 x 54 15 x 44 2003 t 15x²+ 6x + 1 Estept can be seen here given Point (F C2YS 7 = 40 Gola 2 (x - (200 c=600*7 cAnswer the Taylar sories af x 3 - 2 xtL at to (re-2) + 8 (n.2² S (x-23 3 33Kpls like